OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (6, 2) and a directrix of y = 1.

OpenStudy (anonymous):

Deriving equation of the parabola with focus and directrix requires you to use the combined distance formula \[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}=\sqrt{(y_2-y_1)^2}\] Plug in the focus on the left side of the equation and plug in the directrix on the right side.

OpenStudy (anonymous):

would it be 1/2(x-6)^2+3/2 ?

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

Okay so first you substitute your information into the equation. What you do next is to remove the radicals. After you've removed the radicals, distribute the y-term binomials. Simplify and then isolate the x terms. Lastly, isolate the y term to get the equation.

OpenStudy (anonymous):

And yes that's the correct answer, sorry just explaining the process

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

You're welcome!